OpenStudy (anonymous):

Hey Everybody. This is square root problem 2^4+^4=. Can someone help me please? Thanks!

OpenStudy (anonymous):

srqt(2^4+?^4)

OpenStudy (anonymous):

Sorry. These are radicals not square roots. Would the result be the same?

OpenStudy (anonymous):

\[\sqrt{4} +4 \]

OpenStudy (anonymous):

is it something like this

OpenStudy (anonymous):

Smitsa12. How did you draw that symbol. That's the right symbol, but problem is wrong. I can draw it if you show me how to input that symbol. Thank you for your help :) 2^4+^4=

OpenStudy (anonymous):

yes click on equation

OpenStudy (anonymous):

\[2\sqrt{4}+\sqrt{4}=\] This is what it looks like and need help with. Thank you so much :)

OpenStudy (anonymous):

yes both are the same lol

OpenStudy (anonymous):

my equation = urs equation

OpenStudy (anonymous):

Thank you. I never knew how to use the equation box or draw box on here..lol

OpenStudy (anonymous):

\[2\sqrt{4} =\sqrt{4} \]

OpenStudy (anonymous):

and so root over of 4 = howmuch

OpenStudy (anonymous):

Thank you so much..yay!

OpenStudy (anonymous):

that is 2 + 2 = 4

OpenStudy (anonymous):

:P 2 + 2 = 22

OpenStudy (asnaseer):

@JalayiaBarnes1 - if this is your question:\[2\sqrt{4}+\sqrt{4}=\]Then one way to llok at it is by replacing the \(\sqrt{4}\) with, say, "apples" to get:\[2\text{"apples"}+1\text{"apple"}=\]How many "apples" would this result in?

OpenStudy (anonymous):

\[2\sqrt{4} +1\sqrt{4}=3\sqrt{4}=3*2=6\]

OpenStudy (anonymous):

sq rt 4 you know is a number multplied by itself 2*2=4 to get that number hope i am not telling you how to suck eggs as my gran would say

OpenStudy (anonymous):

\[\sqrt{4}+4=2+4=6\]

OpenStudy (anonymous):

Thank you to everyone for your response.